# Dissipation Pathways in a Photosynthetic Complex

### v1

@@ Abstract
Determining how energy flows within and between molecules is crucial for understanding chemical reactions, material properties, and even vital processes such as photosynthesis. While the general principles of energy transfer are well established, elucidating the specific molecular pathways by which energy is funneled remains challenging as it requires tracking energy flow in complex molecular environments. Here, we demonstrate how photon excitation energy is partially dissipated in the light-harvesting Fenna-Matthews-Olson (FMO) complex, mediating the excitation energy transfer from light-harvesting chlorosomes to the photosynthetic reaction center in green sulfur bacteria. Specifically, we isolate the contribution of the protein and specific vibrational modes of the pigment molecules to the energy dynamics. For this, we introduce an efficient computational implementation of a recently proposed theory of dissipation pathways for open quantum systems. Using it and a state-of-the-art FMO model with highly structured and chromophore-specific spectral densities, we demonstrate that energy dissipation is dominated by low-frequency modes ($<$ 800 cm-1) as their energy range is near-resonance with the energy gaps between electronic states of the pigments. We identify the most important mode for dissipation to be in-plane breathing modes ($\sim$200 cm-1) of the bacteriochlorophylls in the complex. Conversely, far-detuned intramolecular vibrations with higher frequencies ($>$ 800 cm-1) play no role in dissipation. Interestingly, the FMO complex first needs to borrow energy from the environment to release excess photonic energy, making the energy dissipation dynamics non-monotonic. Beyond their fundamental value, these insights can guide the development of artificial light-harvesting devices and, more broadly, engineer environments for chemical and quantum control tasks.
Dissipation Pathways|||| Light-harvesting |||| Quantum Dynamics |||| Open Quantum Systems |||| Chemical Design
I Introduction
@@
Photosynthesis is a critically important light-induced process that enables plants, algae, and bacteria to convert solar energy into biochemical fuel with remarkable efficiency.Blankenship (2021); Baker (2008) This transformation begins with the creation of excitons by solar photons, which travel through a number of pigment molecules until reaching the reaction center, where the energy is transformed into chemical fuel.Mirkovic et al. (2017) During this complex energy-transfer process, excitons dissipate excess energy into the nuclear environment. Although the fundamental principles of energy transfer are well established, mapping the precise molecular pathways of this energy flow remains a formidable task as it requires tracking energy dynamics in complex molecular environments.
@@
Specifically, it is not understood how individual components of the photosynthetic system, including the pigment molecules and their surrounding protein environment, contribute to the overall energy dissipation. Understanding these dissipation pathways is essential for characterizing the energy flow and overall dynamics of photosynthetic processes,Mirkovic et al. (2017); Jang and Mennucci (2018); Cao et al. (2020) and is needed to address fundamental questions such as: How do the different components of the pigment molecules contribute to the overall dissipation? Are there specific vibrations that contribute the most to energy dissipation? Is energy transfer a unidirectional process, or do pigment molecules both absorb from and dissipate energy into the environment? Addressing these basic questions requires general strategies to connect chemical structure to energy transfer dynamics in complex molecular systems.
@@
From an experimental perspective, significant spectroscopic efforts have been made to understand the dynamics of energy transfer within photosynthetic complexes.Allodi et al. (2018); Brixner et al. (2005); Engel et al. (2007); Panitchayangkoon et al. (2010); Thyrhaug et al. (2016); Dostál et al. (2016); Duan et al. (2017); Maiuri et al. (2018); Thyrhaug et al. (2018); Higgins et al. (2021a); Segatta et al. (2017) However, it remains challenging to unravel the intricate energy transfer dynamics in complex chemical environments because the spectral congestionArsenault et al. (2020); Lambrev et al. (2020) often prevents disentangling contributions by individual components of the pigments and the protein. Progress has been made through two-dimensional electronic spectroscopy (2DES)Biswas et al. (2022), which allows tracking exciton dynamics along two frequency axes with femtosecond resolution. Nevertheless, even with these advanced techniques, quantifying the contribution of individual nuclear modes to the dissipation remains elusive. This is because the electronic states couple to numerous vibrations, and crucially, 2DES projects this high-dimensional vibronic space onto a reduced measurement dimension, thereby limiting the resolution of individual nuclear contributions to energy dissipation.
@@
From a theory perspective, the task requires a method to decompose energy dissipation into contributions of the protein and specific vibrational modes of the pigment molecules. The challenge is that this requires quantum dynamical information about this complex chemical environment, and the computational cost of obtaining this information remains beyond the reach of state-of-the-art methods in quantum dynamics. For instance, explicit approaches such as the Multi-configuration time-dependent Hartree (MCTDH)Beck et al. (2000) accurately model environmental dynamics through direct wavefunction propagation, yet they become computationally intractable for macroscopic chemical environments. By contrast, quantum master equations and related techniquesKundu et al. (2022); Strathearn et al. (2018); Varvelo et al. (2021); Bose and Walters (2022) can treat complex chemical environments implicitly by focusing on the effect in the system’s dynamics. This scalability, however, comes at the cost of discarding information about the environment’s quantum dynamics.
@@
In response to this challenge, we recently developedKim and Franco (2024a, 2021) a general quantum theory to quantify and resolve dissipation pathways of quantum systems immersed in highly structured quantum thermal environments. The theory, based on second-order perturbation theory of the off-diagonal system’s couplings, captures the energy dissipated into individual environmental modes while avoiding explicitly propagating the quantum dynamics of the environment, thereby opening a systematic and computationally tractable path to investigate the energy flow between the system and its environment. Here, we introduce an efficient computational implementation (see Theoretical Methods) of this theory that now enables access to the dissipation pathways of complex molecular systems while ensuring that the computational cost remains tractable.
@@
Here, we use this strategy to elucidate how photon excitation energy is dissipated in the Fenna-Matthews-Olson (FMO) complexFenna and Matthews (1975). This complex mediates the transfer of excitation energy from light-harvesting chlorosomes to the photosynthetic reaction center in green sulfur bacteria.Blankenship (2021); Jang and Mennucci (2018) While the structure and exciton dynamics in the FMO have been well studiedAdolphs and Renger (2006); González-Soria et al. (2020); Kell et al. (2016); Christensson et al. (2012); Gillis and Jones (2015); Rivera et al. (2013); Wendling et al. (2000); Lee and Coker (2016); Brixner et al. (2005); Engel et al. (2007); Kim et al. (2018); Higgins et al. (2021a); Orf et al. (2016); Malý et al. (2016); Abramavicius and Abramavicius (2014); Padula et al. (2017); Schulze and Kühn (2015); Olbrich et al. (2011a); Moix et al. (2011); Chandrasekaran et al. (2015); Milder et al. (2010); Tronrud et al. (2009); Schmidt am Busch et al. (2011); Camara-Artigas et al. (2003); Olbrich et al. (2011b); Chen et al. (2022), the dissipation pathways in the complex have remained elusive to both theory and experiment. Specifically, we isolate the contributions of the protein and intramolecular vibrational modes of the pigment molecules to the energy dynamics. We utilize highly structured and pigment-specific spectral densities to characterize the frequencies of protein and intramolecular vibrations, as well as their coupling strengths to electronic excitations in the pigments.Kim et al. (2018) Since these highly structured spectral densities represent high-dimensional environments with $20+$ vibrational features per pigment, computing the dissipation pathways is beyond the reach of any other state-of-the-art method in quantum dynamics.
@@
As discussed below, vibrations with frequencies $<$ 800 cm-1 dominate the energy transfer dynamics, with in-plane breathing modes of the bacteriochlorophylls being the most important ones. Interestingly, the simulations also show that the FMO complex dissipates excess photonic energy through a non-monotonic process, as it initially draws energy from the thermal environment before onsetting the overall dissipative process.
@@
The introduced strategy provides a path to disentangle the contribution of individual molecules and nuclear vibrations to the energy dynamics during photosynthetic events.Mirkovic et al. (2017); Jang and Mennucci (2018); Cao et al. (2020) This is important for establishing the connection between molecular structure and energy dynamics. Beyond its fundamental value, this connection can be helpful for the design of artificial light-harvesting systems,Hart et al. (2021); Wasielewski (2009); Tang et al. (2024) and, more broadly, to engineer environments for chemical and quantum control tasks.Kienzler et al. (2014); Hsu et al. (2017); Campos-Gonzalez-Angulo et al. (2019); Ng et al. (2020); Gustin et al. (2025a); Hu et al. (2022)
@@
Our efforts augment and complement previous efforts to investigate energy pathways in molecular arrays. In particular, there are important strategies to track the energy flow in proteins through vibrational energy transfers using atomisticSagnella et al. (2000); Bu and Straub (2003) and coarse-grainedIshikura et al. (2015); Xu and Leitner (2014a, b); Leitner et al. (2015); Leitner (2009); Agbo et al. (2015) classical molecular dynamics simulations. Our approach provides a fully quantum perspective that focuses on the quantification of the dissipation of electronic excitation into nuclear modes. Our efforts also complement strategies developed by Bittner and coworkersYang et al. (2017); Pereverzev et al. (2009) to isolate the most important coordinate in complex electron transfer processes using electronic structure methods. Such models can be incorporated into our theoretical framework, enabling the analysis of their associated dissipation pathways.
Dissipation Pathways in the FMO Complex
@@
We analyze how energy is dissipated into the thermal environment in the Fenna-Matthews-Olson (FMO) complex, essential for efficient light harvesting in green sulfur bacteria such as C. tepidum and P. aestuarii. The FMO complex acts as a molecular bridge, facilitating the transfer of excitation energy from the chlorosome, where photons are absorbed, to the reaction center, where this energy is converted into chemical fuel. While this transfer is mediated by exciton dynamics within a trimeric FMO complex, research suggests that each exciton’s pathway is confined to a single monomer during transferJang and Mennucci (2018), each housing seven bacteriochlorophyll (Bchl) molecules (see Fig. 1a). Therefore, we focus on a single monomer.
@@
To describe the energies and couplings of the excited electronic states in the 7-Bchl array, we adopt the Adolphs and RengerAdolphs and Renger (2006) Hamiltonian (see Sec. SI in the Supplementary Information) as it accurately reproduces the optical spectra of the FMO complex. In turn, the interaction between the Bchls and the surrounding environment is described by Bchl-specific spectral densities, computed by Kim and coworkersKim et al. (2018), which accurately capture the nuclear environment’s vibrational frequencies ($\omega$) and their coupling strengths with the Bchls electronic excited state. These spectral densities consist of broad features at frequencies below $200$ cm-1 that account for protein contributions and a series of sharp peaks to describe the influence of intramolecular vibrational modes. In Fig. 1b, we detail a segment of the spectral density scaled by the environment frequency, $J(\omega)/\omega$, for the chromophores that are most important in the process of absorption/dissipation of energy. The remaining spectral densities are presented in Fig. S1.
@@
Our analysis is based on a recently proposed quantum theory of dissipation pathwaysKim and Franco (2024a, b), see Theoretical Methods for details, which is used to quantify the amount of energy channeled from the Bchls in the FMO into the complex’s vibrational modes, including both protein and intramolecular contributions. The applicability of our theory, which assumes weak electronic coupling, to compute dissipation pathways in the FMO complex is validated against the numerically exact hierarchical equation of motion (HEOM)Tanimura (2020); Ikeda and Scholes (2020); Kim (2022) using illustrative models (see Secs. SII-SIII in the Supplementary Information), showing semi-quantitative agreement. We note that direct HEOM calculations of dissipation pathways incorporating the highly structured spectral densities of the FMO complex are currently computationally prohibitive. Thus, our benchmark, using representative models that capture the essential features of the FMO, represents the state of the art.
@@
Figure 1 shows the contributions of specific environmental modes at frequency $\omega$ to the overall asymptotic energy absorption/dissipation in the overall complex and in each chromophore when the initial excitation is placed in Bchl1, the site closest to the chlorosome.Wen et al. (2009) This electronic excitation moves through the complex until it reaches Bchl3 and Bchl4, the chromophores closest to the reaction center.Adolphs and Renger (2006) In Fig. 1a, chromophores are labeled with bold numbers, and the intensity of the green (or red) color indicates the relative amount of energy dissipated (or absorbed) at each chromophore in the overall process.
@@
Figure 1c shows the overall frequency-resolved dissipation profile, which reveals that the dissipation of energy is driven by the low-frequency modes ($<$ 800 cm-1). This is because their energy is near-resonant with the energy gaps between excited electronic states in the complex, see Fig. 1d. In particular, we identify environmental modes around 200 cm-1 as the most important modes for dissipation in the FMO complex, giving the most important contribution in Fig. 1c. Remarkably, despite being present in the spectral densitiesKim et al. (2018), see Fig. S1, high-frequency vibrational modes ($>$ 800 cm-1) do not contribute to the energy dissipation in this complex as their energy is far detuned from the energy gaps between excited electronic states. This is consistent with previous analyses in the FMO complexAbramavicius and Abramavicius (2014); Malý et al. (2016); Olbrich et al. (2011a); Schulze and Kühn (2015); Padula et al. (2017); Wu et al. (2012) that showed that high-frequency components are not necessary to capture exciton dynamics accurately.
@@
Vibrational Mode Driving Dissipation
To elucidate the vibrational modes around 200 cm-1 that predominantly contribute to dissipation in the FMO complex across all chromophores, we developed a computational strategy to extract the representative vibrational motion from the molecular dynamics trajectories used for constructing the spectral densities (see Sec. SIV in the Supplementary Information for details). Figure 2 shows the isolated vibrational mode around 200 cm-1 for Bchl3, displaying only the Bchl core for enhanced clarity. Interestingly, this vibrational mode, most responsible for dissipation, exhibits a breathing motion pattern and is confined to the plane with no out-of-plane components. We do not show other chromophores as the modes around 200 cm-1 exhibit remarkable similarity across all chromophores, including Bchl1, which absorbs energy from the environment. A set of animations showing the vibrational mode for every Bchl around 200 cm-1 are available in the supplementary materials. Importantly, the relevance of this vibrational frequency around 200 cm-1 has been previously noted in 2DES experiments in the FMO complex where including this vibrational mode in the spectral density was essential for obtaining energy transfer rates that align with experimental observations.Higgins et al. (2021a)
@@
Time-Dependent Dissipation
We now go beyond the overall dissipation, and scrutinize the role of each Bchl during the energy transfer dynamics. Fig. 3 shows the cumulative dissipation over time for the four most significant chromophores and for the overall complex. Fig. 3a shows that, initially, the dissipation is negative, indicating that the system is absorbing energy from the environment to facilitate uphill energy transfer from Bchl1 to Bchl2 (see energy level diagram in Fig. 3d). Interestingly, both Bchl1 and Bchl2 collectively absorb energy during this phase. This initial energy absorption is consistent with the uphill energy transfer signals involving Bchl1 and Bchl2 observed experimentally in two-dimensional electronic spectroscopy (2DES) experiments measured at 40 femtoseconds (fs).Higgins et al. (2021b)
@@
This analysis shows that energy flow within the FMO complex is not a monotonic process. Some chromophores absorb energy from or dissipate energy into the environment throughout the dynamics, while others can absorb or dissipate energy at different time scales. Furthermore, the analysis shows that not all vibrational modes are equally important for the energy transfer dynamics in the FMO complex. In fact, we identified the in-plane breathing modes around 200 cm-1 as the most important for energy transfer.
Conclusions
@@
In conclusion, for the first time, we have successfully isolated the dissipation pathways in the Fenna-Matthews-Olson complex by quantifying contributions by individual environmental components to the energy dynamics. This was achieved through an efficient computational implementation of our recently developed quantum theory of dissipation pathwaysKim and Franco (2024a, b) combined with a state-of-the-art FMO model incorporating highly structured and Bchl-specific spectral densities.
@@
Our analysis revealed that energy dissipation in the FMO complex is predominantly driven by low-frequency vibrational modes ($<$ 800 cm-1), with in-plane breathing modes around 200 cm-1 being the most dominant. High-frequency modes ($>$ 800 cm-1) play a negligible role despite their presence in the spectral densities as they are far detuned from the excited state electronic energy gaps in the FMO complex. Furthermore, we demonstrated that energy flow within the FMO complex is non-monotonic. Specifically, to release excess photonic energy, the FMO complex must first absorb energy from the environment. We identify BChl1, closest to the chlorosome, as the primary mediator of initial energy absorption and BChl3, nearest to the reaction center, as the primary mediator of energy dissipation.

### v2

@@ Abstract
Determining how energy flows within and between molecules is crucial for understanding chemical reactions, material properties, and even vital processes such as photosynthesis. While the general principles of energy transfer are well established, elucidating the specific molecular pathways by which energy is funneled remains challenging as it requires tracking energy flow in complex molecular environments. Here, we demonstrate how photon excitation energy is partially dissipated in the light-harvesting Fenna-Matthews-Olson (FMO) complex, mediating the excitation energy transfer from light-harvesting chlorosomes to the photosynthetic reaction center in green sulfur bacteria. Specifically, we isolate the contribution of the protein and specific vibrational modes of the pigment molecules to the energy dynamics. For this, we introduce an efficient computational implementation of a recently proposed theory of dissipation pathways for open quantum systems, based on second-order perturbation theory in the electronic couplings. Using it and a state-of-the-art FMO model with highly structured and chromophore-specific spectral densities, we demonstrate that energy dissipation is dominated by low-frequency modes (<< 800 cm-1) as their energy range is near-resonance with the energy gaps between electronic states of the pigments. We identify the most important mode for dissipation to be in-plane breathing modes ($\sim$200 cm-1) of the bacteriochlorophylls in the complex. Conversely, far-detuned intramolecular vibrations with higher frequencies (>> 800 cm-1) play no role in dissipation. Interestingly, the FMO complex first needs to borrow energy from the environment to release excess photonic energy, indicating that the energy exchange between the system and the thermal environment is not strictly unidirectional in time but involves a transient thermally activated step. Beyond their fundamental value, these insights can guide the development of artificial light-harvesting devices and, more broadly, engineer environments for chemical and quantum control tasks.
keywords:
American Chemical Society,
[…]
{tocentry}[Uncaptioned image]
@@
Photosynthesis is a critically important light-induced process that enables plants, algae, and bacteria to convert solar energy into biochemical fuel with remarkable efficiency.blankenship2021molecular, baker2008chlorophyll This transformation begins with the creation of excitons by solar photons, which travel through a number of pigment molecules until reaching the reaction center, where the energy is transformed into chemical fuel.mirkovic2017light During this complex energy-transfer process, excitons dissipate excess energy into the nuclear environment. Although the fundamental principles of energy transfer are well established, mapping the precise molecular pathways of this energy flow remains a formidable task as it requires tracking energy dynamics in complex molecular environments.
@@
Specifically, it is not understood how individual components of the photosynthetic system, including the pigment molecules and their surrounding protein environment, contribute to the overall energy dissipation. Understanding these dissipation pathways is essential for characterizing the energy flow and overall dynamics of photosynthetic processes,mirkovic2017light, jang2018delocalized, cao2020quantum and is needed to address fundamental questions such as: How do the different components of the pigment molecules contribute to the overall dissipation? Are there specific vibrations that contribute the most to energy dissipation? Is energy transfer a unidirectional process, or do pigment molecules both absorb from and dissipate energy into the environment? Addressing these basic questions requires general strategies to connect chemical structure to energy transfer dynamics in complex molecular systems.
@@
From an experimental perspective, significant spectroscopic efforts have been made to understand the dynamics of energy transfer within photosynthetic complexes.allodi2018redox, brixner2005two, engel2007evidence, panitchayangkoon2010long, thyrhaug2016exciton, dostal2016situ, duan2017nature, maiuri2018coherent, thyrhaug2018identification, higgins2021photosynthesis, segatta_quantum_2017 However, it remains challenging to unravel the intricate energy transfer dynamics in complex chemical environments because the spectral congestionarsenault2020role, lambrev2020insights often prevents disentangling contributions by individual components of the pigments and the protein. Progress has been made through two-dimensional electronic spectroscopy (2DES)biswas2022coherent, which allows tracking exciton dynamics along two frequency axes with femtosecond resolution. Nevertheless, even with these advanced techniques, quantifying the contribution of individual nuclear modes to the dissipation remains elusive. This is because the electronic states couple to numerous vibrations, and crucially, 2DES projects this high-dimensional vibronic space onto a reduced measurement dimension, thereby limiting the resolution of individual nuclear contributions to energy dissipation.
@@
From a theory perspective, the task requires a method to decompose energy dissipation into contributions of the protein and specific vibrational modes of the pigment molecules. The challenge is that this requires quantum dynamical information about this complex chemical environment, and the computational cost of obtaining this information remains beyond the reach of state-of-the-art methods in quantum dynamics. For instance, explicit approaches such as the Multi-configuration time-dependent Hartree (MCTDH)beck2000multiconfiguration accurately model environmental dynamics through direct wavefunction propagation, yet they become computationally intractable for macroscopic chemical environments. Trajectory-based mixed quantum-classical methods can also treat complex chemical environments.sokolov2024non, kim2012all However, they suffer from artifacts like zero-point energy leakage and negative populations that lead to unphysical dissipation computations.kim2014improving, kim2020toward Moreover, their classical treatment of the environment does not correctly capture important quantum processes such as spontaneous emission. By contrast, quantum master equations and related techniqueskundu_tight_2022, tempo, varvelo2021formally, Bose2022multisite can treat complex chemical environments implicitly by focusing on the effect in the system’s dynamics. This scalability, however, comes at the cost of discarding information about the environment’s quantum dynamics.
@@
In response to this challenge, we recently developedkim2024general1, kim2021theory a general fully quantum theory to quantify and resolve dissipation pathways of quantum systems immersed in highly structured quantum thermal environments, including both harmonic and anharmonic baths. The theory, based on second-order perturbation theory of the off-diagonal system’s couplings, captures the energy dissipated into individual environmental modes while avoiding explicitly propagating the quantum dynamics of the environment, thereby opening a systematic and computationally tractable path to investigate the energy flow between the system and its environment. Here, we introduce an efficient computational implementation (see Theoretical Methods) of this theory that now enables access to the dissipation pathways of complex molecular systems while ensuring that the computational cost remains tractable.
@@
We use this strategy to elucidate how photon excitation energy is dissipated in the Fenna-Matthews-Olson (FMO) complexfenna1975chlorophyll. This complex mediates the transfer of excitation energy from light-harvesting chlorosomes to the photosynthetic reaction center in green sulfur bacteria.blankenship2021molecular, jang2018delocalized While the structure and exciton dynamics in the FMO have been well studiedadolphs2006proteins, gonzalez-soria_parametric_2020, kell_effect_2016, christensson_origin_2012, gillis_theoretical_2015, rivera_influence_2013, wendling_electronvibrational_2000, lee_modeling_2016, brixner2005two, engel2007evidence, kim2018excited, higgins2021photosynthesis, orf_evidence_2016, maly_role_2016, abramavicius_excitation_2014, padula_chromophore-dependent_2017, schulze_explicit_2015, olbrich_theory_2011, moix2011efficient, chandrasekaran2015influence, milder2010revisiting, tronrud2009structural, schmidt2011eighth, camara2003structure, olbrich2011atomistic, chen2022simulation, the dissipation pathways in the complex have remained elusive to both theory and experiment. Specifically, we isolate the contributions of the protein and intramolecular vibrational modes of the pigment molecules to the energy dynamics. We utilize highly structured and pigment-specific spectral densities to characterize the frequencies of protein and intramolecular vibrations, as well as their coupling strengths to electronic excitations in the pigments.kim2018excited Since these highly structured spectral densities represent high-dimensional environments with $20+$ vibrational features per pigment, computing the dissipation pathways is beyond the reach of any other state-of-the-art method in quantum dynamics.
@@
As discussed below, vibrations with frequencies << 800 cm-1 dominate the energy transfer dynamics, with in-plane breathing modes of the bacteriochlorophylls being the most important ones. While the importance of low-frequency vibrations in the FMO complex has been recognizednalbach2015vibronically our work moves beyond this, elucidating how and why specific modes govern the dissipation dynamics with no a priori assumptions. Interestingly, the simulations also show that the FMO complex dissipates excess photonic energy through a non-monotonic process, as it initially draws energy from the thermal environment before onsetting the overall dissipative process.
@@
The introduced strategy provides a path to disentangle the contribution of individual molecules and nuclear vibrations to the energy dynamics during photosynthetic events.mirkovic2017light, jang2018delocalized, cao2020quantum This is important for establishing the connection between molecular structure and energy dynamics. Beyond its fundamental value, this connection can be helpful for the design of artificial light-harvesting systems,hart2021engineering, wasielewski2009self, tang2024simulating and, more broadly, to engineer environments for chemical and quantum control tasks.Kienzler2014, Hsu2017, CamposGonzalezAngulo2019, Ng2020, gustin_decoherence_2025, hu2022tuning
@@
Our efforts augment and complement previous efforts to investigate energy pathways in molecular arrays. In particular, there are important strategies to track the energy flow in proteins through vibrational energy transfers using atomisticsagnella2000time, bu2003simulating and coarse-grainedishikura2015energy, xu2014vibrational, xu2014communication, leitner2015vibrational, leitner2009frequency, agbo2015vibrational classical molecular dynamics simulations. Our approach provides a fully quantum perspective that focuses on the quantification of the dissipation of electronic excitation into nuclear modes. Our efforts also complement strategiesyang2017identifying, pereverzev2009energy developed to isolate the most important coordinate in complex electron transfer processes using electronic structure methods. Such models can be incorporated into our theoretical framework, enabling the analysis of their associated dissipation pathways.
@@
We analyze how energy is dissipated into the thermal environment in the Fenna-Matthews-Olson (FMO) complex, essential for efficient light harvesting in green sulfur bacteria such as C. tepidum and P. aestuarii. The FMO complex acts as a molecular bridge, facilitating the transfer of excitation energy from the chlorosome, where photons are absorbed, to the reaction center, where this energy is converted into chemical fuel. While this transfer is mediated by exciton dynamics within a trimeric FMO complex, research suggests that each exciton’s pathway is confined to a single monomer during transferjang2018delocalized, each housing eight bacteriochlorophyll (Bchl) molecules (see Fig. 1a). Therefore, we focus on a single monomer.
@@
To describe the energies and couplings of the excited electronic states in the Bchl array, we adopt the Adolphs and Rengeradolphs2006proteins Hamiltonian (see Sec. SI in the Supplementary Information). The model includes 7 of the 8 BChl as one of them acts as an incoherent energy funnel exlusively connected to BChl1.moix2011efficient While reported site energies can differ by up to 100 cm-1 across studieshigashi2016quantitative, the overall energy landscape and transfer pathways remain consistent across different parameter sets. The Adolphs and Ranger parameterization provides a well-established model that captures this consistent landscape and accurately reproduces the optical spectra of the FMO complex.adolphs2006proteins The interaction between the Bchls and the surrounding environment is described by Bchl-specific spectral densities, computed by Kim and coworkerskim2018excited, which accurately capture the nuclear environment’s vibrational frequencies ($\omega$) and their coupling strengths with the Bchls electronic excited state with a frequency resolution of $10^{-4}$ cm-1. These spectral densities show good agreement with available experiments and consist of broad features at frequencies below $200$ cm-1 that account for protein contributions and a series of sharp peaks to describe the influence of intramolecular vibrational modes. In Fig. 1b, we detail a segment of the spectral density scaled by the environment frequency, $J(\omega)/\omega$, for the chromophores that are most important in the process of absorption/dissipation of energy. The remaining spectral densities are presented in Fig. S1. The reliability of our spectral densities, particularly in the challenging low-frequency region, is ensured by the extensive 100 ns MD trajectory employed in its construction. This approach provides a high frequency resolution of 3 $\times$ 10-4 cm-1.kim2018excited
@@
Our analysis is based on a recently proposed fully quantum theory of dissipation pathwayskim2024general1, kim2024general2, see Theoretical Methods for details, which is used to quantify the amount of energy channeled from the Bchls in the FMO into the complex’s vibrational modes, including both protein and intramolecular contributions. The applicability of our theory, which assumes weak electronic coupling, to compute dissipation pathways in the FMO complex is validated against the numerically exact hierarchical equation of motion (HEOM)tanimura2020numerically, Ikeda2020, kim_extracting_2022 using illustrative models (see Secs. SII-SIII in the Supplementary Information), showing semi-quantitative agreement. We note that direct HEOM calculations of dissipation pathways incorporating the highly structured spectral densities of the FMO complex are currently computationally prohibitive. Thus, our benchmark, using representative models that capture the essential features of the FMO, represents the state of the art.
@@
Figure 1 shows the contributions of specific environmental modes at frequency $\omega$ to the overall asymptotic energy absorption/dissipation in the overall complex and in each chromophore when the initial excitation is placed in Bchl1, the site closest to the chlorosome.wen2009membrane This electronic excitation moves through the complex until it reaches Bchl3 and Bchl4, the chromophores closest to the reaction center.adolphs2006proteins In Fig. 1a, chromophores are labeled with bold numbers, and the intensity of the green (or red) color indicates the relative amount of energy dissipated (or absorbed) at each chromophore in the overall process.
@@
Figure 1c shows the overall frequency-resolved dissipation profile, which reveals that the dissipation of energy is driven by the low-frequency modes (<< 800 cm-1). This is because their energy is near-resonant with the energy gaps between excited electronic states in the complex, see Fig. 1d. In particular, we identify environmental modes around 200 cm-1 as the most important modes for dissipation in the FMO complex, giving the most important contribution in Fig. 1c. Remarkably, despite being present in the spectral densitieskim2018excited, see Fig. S1, high-frequency vibrational modes (>> 800 cm-1) do not contribute to the energy dissipation in this complex as their energy is far detuned from the energy gaps between excited electronic states. This is consistent with previous analyses in the FMO complexabramavicius_excitation_2014, maly_role_2016, olbrich_theory_2011, schulze_explicit_2015, padula_chromophore-dependent_2017, wu_efficient_2012 that showed that high-frequency components are not necessary to capture exciton dynamics accurately.
@@
To elucidate the vibrational modes around 200 cm-1 that predominantly contribute to dissipation in the FMO complex across all chromophores, we developed a computational strategy to extract the representative vibrational motion from the molecular dynamics trajectories used for constructing the spectral densities (see Sec. SIV in the Supplementary Information for details). Figure 2 shows the isolated vibrational mode around 200 cm-1 for Bchl3, displaying only the Bchl core for enhanced clarity. Interestingly, this vibrational mode, most responsible for dissipation, exhibits a breathing motion pattern and is confined to the plane with no out-of-plane components. We do not show other chromophores as the modes around 200 cm-1 exhibit remarkable similarity across all chromophores, including Bchl1, which absorbs energy from the environment. A set of animations showing the vibrational mode for every Bchl around 200 cm-1 are available in the supplementary materials. Importantly, the relevance of this vibrational frequency around 200 cm-1 has been previously noted in 2DES experiments in the FMO complex where including this vibrational mode in the spectral density was essential for obtaining energy transfer rates that align with experimental observations.higgins2021photosynthesis Furthermore, theoretical investigationsnalbach2015vibronicallyhave also found it necessary to include this mode on an ad hoc basis to reproduce experimental results, reinforcing the consensus of its critical role. By contrast, our work does not presuppose its importance. Instead, the dominant role of the 200 cm-1 vibration is a direct outcome of the system’s dynamics, providing a first-principles validation of its function and revealing the underlying molecular mechanism driving the dissipation.
To ensure that these findings are not dependent on the initial excitation site, we conducted an identical simulation with the excitation initiated on BChl6. These calculations, detailed in Sec. SV in the Supplementary Information, yield a consistent picture of the energy dissipation dynamics. The process remains dominated by low-frequency nuclear modes (<< 800 cm-1), with the vibrational mode near 200 cm-1 consistently emerging as the most significant contributor to energy relaxation. This confirms that the identified dissipation channel is a robust feature of the system’s dynamics.
@@
We now go beyond the overall dissipation, and scrutinize the role of each Bchl during the energy transfer dynamics. Fig. 3 shows the cumulative dissipation over time for the four most significant chromophores and for the overall complex. Fig. 3a shows that, initially, the dissipation is negative, indicating that the system is absorbing energy from the environment to facilitate uphill energy transfer from Bchl1 to Bchl2 (see energy level diagram in Fig. 3d). Interestingly, both Bchl1 and Bchl2 collectively absorb energy during this phase. This initial energy absorption is consistent with the uphill energy transfer signals involving Bchl1 and Bchl2 observed experimentally in two-dimensional electronic spectroscopy (2DES) experiments measured at 40 femtoseconds (fs).higgins2021redox
@@
This analysis shows that energy flow within the FMO complex is not unidirectional. Some chromophores absorb energy from or dissipate energy into the environment throughout the dynamics, while others can absorb or dissipate energy at different time scales. Furthermore, the analysis shows that not all vibrational modes are equally important for the energy transfer dynamics in the FMO complex. In fact, we identified the in-plane breathing modes around 200 cm-1 as the most important for energy transfer.
@@
In conclusion, for the first time, we have successfully isolated the dissipation pathways in the Fenna-Matthews-Olson complex by quantifying contributions by individual environmental components to the energy dynamics. This was achieved through an efficient computational implementation of our recently developed quantum theory of dissipation pathwayskim2024general1, kim2024general2 combined with a state-of-the-art FMO model incorporating highly structured and Bchl-specific spectral densities.
@@
Our analysis revealed that energy dissipation in the FMO complex is predominantly driven by low-frequency vibrational modes (<< 800 cm-1), with in-plane breathing modes around 200 cm-1 being the most dominant. High-frequency modes (>> 800 cm-1) play a negligible role despite their presence in the spectral densities as they are far detuned from the excited state electronic energy gaps in the FMO complex. Furthermore, we demonstrated that energy flow within the FMO complex is non-monotonic, as it involves a transient back-and-forth energy exchange, reflecting a thermally activated step prior to overall dissipation. Specifically, to release excess photonic energy, the FMO complex must first absorb energy from the environment. We identify BChl1, closest to the chlorosome, as the primary mediator of initial energy absorption and BChl3, nearest to the reaction center, as the primary mediator of energy dissipation.